amat:

(08.07) Find the 10th partial sum of the summation of negative 5i plus 2, from i equals 1 to infinity. (1 point) 222 240 255 −255

yourdootcome:

the sum of -5i + 2 from i = 1 to 10 is equal to -5(sum of i) from 1 to 10 + (sum of 2) from 1 to 10 then that equals: - 5(1 + 10) * 10/2 + 2(10) which simplified equals: -5 * 11 * 5 + 20 = -275 + 20 = -255